OpenStudy (anonymous):

giving medals! The second side of a triangle is 3 more meters than the first side. The third side is twice more than the first side. If the perimeter of the triangle is 47 meters, find the length of the second side.

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (kaliyah0524):

Let the first side be triangle be x Then second side of a triangle = x+3 and accordingly third side = 2x Perimeter of the triangle = sum of all the sides = 47 x+x+3+2x = 47 4x+3 = 47 Subtracting 3 from either side results in.. 4x+3-3 = 47-3 4x= 44 Dividing both sides by 4 x=11 So second side of a triangle = x+3 = 11+3 Ans : 14 meter

OpenStudy (anonymous):

thanks you sure right :)

OpenStudy (kaliyah0524):

Glad to help (: